To notify the workers on a farm that dinner is ready, a cook rings a triangle producing sound waves of frequency 1024 Hz. If it takes the sound 0.55 s to reach the farm workers, 190 m away, calculate the wavelength of the sound wave.


1
Expert's answer
2021-01-26T19:26:39-0500

Let's first find the speed of the sound wave:


"v=\dfrac{s}{t}=\dfrac{190\ m}{0.55\ s}=345.45\ \dfrac{m}{s}."

Finally, from the wave speed formula we can find the wavelength of the sound wave:


"v=f\lambda,""\lambda=\dfrac{v}{f}=\dfrac{345.45\ \dfrac{m}{s}}{1024\ Hz}=0.34\ m."